# Balanced Token Pruning: Accelerating Vision Language Models Beyond Local Optimization

### v1

@@ Abstract
Large Vision-Language Models (LVLMs) have shown impressive performance across multi-modal tasks by encoding images into thousands of tokens. However, the large number of image tokens results in significant computational overhead, and the use of dynamic high-resolution inputs further increases this burden. Previous approaches have attempted to reduce the number of image tokens through token pruning, typically by selecting tokens based on attention scores or image token diversity. Through empirical studies, we observe that existing methods often overlook the joint impact of pruning on both the current layer’s output (local) and the outputs of subsequent layers (global), leading to suboptimal pruning decisions. To address this challenge, we propose Balanced Token Pruning (BTP), a plug-and-play method for pruning vision tokens. Specifically, our method utilizes a small calibration set to divide the pruning process into multiple stages. In the early stages, our method emphasizes the impact of pruning on subsequent layers, whereas in the deeper stages, the focus shifts toward preserving the consistency of local outputs. Extensive experiments across various LVLMs demonstrate the broad effectiveness of our approach on multiple benchmarks. Our method achieves a 78% compression rate while preserving 96.7% of the original models’ performance on average.
@@ 1 Introduction
Recent advances in Large Vision-Language Models (LVLMs) chen2024internvl ; gao2024mini ; liu2023improvedllava ; liu2023llava ; wang2024qwen2 have substantially improved visual understanding. These models typically employ a visual encoder to convert images into discrete tokens, which are then processed jointly with textual tokens by a large language model backbone. The incorporation of visual information significantly increases the total number of input tokens bai2025qwen2 ; liu2024llavanext , a problem further amplified when handling high-resolution images. In scenarios with constrained computational resources and limited memory capacity, the large volume of visual tokens leads to considerable increases in inference latency and memory usage, thereby limiting the practical applicability of LVLMs in resource-limited environments.
Prior studies have demonstrated that visual tokens often exhibit significant redundancy chen2024image ; lin2025boosting . Consequently, visual token pruning has been proposed as an effective strategy to reduce input redundancy and enhance computational efficiency yang2024visionzip ; shang2024llava ; huang2024dynamic ; zhang2025llava ; ye2024atp . Visual token pruning faces two fundamental challenges: identifying the most important visual tokens and determining the appropriate layers for pruning. Existing token pruning strategies can be broadly classified into two categories: attention-based methods that leverage text-image interactions chen2024image ; xing2024pyramiddrop , and diversity-based methods that exploit the heterogeneity of visual representations alvar2025divprune . However, the impact of their distinct optimization objectives on overall model performance remains underexplored, and a systematic comparison between them is largely absent. Moreover, when it comes to pruning layer selection, existing methods rely heavily on validation performance and manually defined settings, lacking principled guidance based on the model’s intrinsic properties.
@@
To address these problems, we first explore the nature of image token pruning from an intuitive perspective: its impact on the current layer’s (local) output and its influence on the outputs of subsequent pruning layers (global). We begin by visualizing the spatial distribution of image tokens that receive higher attention from text tokens across different layers. As shown in Figure 1, we observe that the image tokens attended by text tokens vary across different layers. This indicates that pruning solely based on the current layer tends to overlook its impact on subsequent layers. Then we further compare the similarity between the hidden state of the last word in the input prompt under two pruning methods and that of the original model.
@@ 2.1 Large Vision-Language Models
Recent progress in large vision language models (LVLMs) has been substantially accelerated by the open-sourcing of foundation models like LLaMA touvron2023llama and Vicuna zheng2023judging . Representative models, including LLaVA liu2023improvedllava ; liu2024llavanext ; liu2023llava , Qwen-VL bai2025qwen2 ; wang2024qwen2 , and InternVL chen2024internvl ; gao2024mini leverage vision encoders radford2021learningtransferablevisualmodels ; li2022blipbootstrappinglanguageimagepretraining ; choraria2024semanticallygroundedqformerefficient to encode images into visual tokens, which are then integrated into the language model for unified multimodal representation and understanding. For example, LLaVA-1.5 encodes image into 576 visual tokens using a single-scale encoder. As these models increasingly support high-resolution visual inputs bai2025qwen2 ; liu2024llavanext , the number of visual tokens grows. Using a multi-resolution encoding strategy, LLaVA-NeXT can generate up to 2,880 tokens per image.
@@ 2.2 Visual Token Pruning
Early efforts to reduce visual token redundancy primarily focus on attention-based pruning chen2024efficient ; huang2025dynamicllavaefficientmultimodallarge ; zhang2025llavaminiefficientimagevideo . For example, FastV chen2024image prunes visual tokens with low attention scores after the filtering layer, with subsequent layers processing only the remaining token. Another approach, VTW lin2025boosting , adopts a complete token elimination strategy, removing all visual tokens after a specified layer. PyramidDrop xing2024pyramiddrop introduces a more sophisticated approach, performing stage-wise pruning throughout the transformer, ranking visual tokens by their attention scores to the instruction token at each stage and progressively discarding the least informative ones. Compared to attention-based methods, diversity-based methods prioritize retaining a richer variety of semantic information. For instance, DivPrune alvar2025divprune formulate token pruning as a Max-Min Diversity Problem porumbel2011simple ; resende2010grasp . Additionally, some methods fuse remaining tokens into retained tokens through token fusion such as LLaVA-PruMerge shang2024llava and VisionZip yang2024visionzip . Different from prior methods, our method jointly considers the impact of pruning on both the current layer and subsequent layers.
@@ 3.1 Visual token processing
In the prefilling stage, images and texts are first encoded into embedding vectors (tokens), which are then processed by LVLM. We denote the input token sequence as $\mathbf{X}$ which consists of the system prompt $\mathbf{X}_{S}$, the image tokens $\mathbf{X}_{I}$ and text tokens $\mathbf{X}_{T}$, $\mathbf{X}$ = ($\mathbf{X}_{S}$,$\mathbf{X}_{I}$,$\mathbf{X}_{T}$) . $\mathbf{X}$ is then fed into the LLM backbone composed of N decoder layers. For the l-th decoder layer, we denote the input as $\mathbf{X}^{(l)}$ and the layer output $\mathbf{X}^{(l+1)}$ is:
@@ 4.1 Limitations of existing methods
We analyze the impact of pruning image tokens on the subsequent text and response tokens. From Equations 1 and 2, we can see that pruning image tokens at $l$-th layer mainly affects the layer output $\mathbf{X}^{(l+1)}$ by changing the attention output, which is a weighted sum of the value vectors $V_{l}$. If the norms of the $V_{l}$ are similar, selecting image tokens with high importance scores defined in 3 effectively reduces the difference between the layer output before and after pruning. We provide supporting evidence for this assumption in the Appendix A.1. Formally, given original $l$-th layer output $\mathbf{X}_{origin}^{(l+1)}$ and pruned $l$-th layer output $\mathbf{X}^{(l+1)}_{pruned}$ , distance metric function $D(\cdot,\cdot)$, we can define the objective function $\mathcal{L}_{atten}$ of attention-based methods chen2024image ; xing2024pyramiddrop as:
@@ 4.2 Balanced token pruning with joint local and global objectives
Based on the above analysis, we argue that an effective token pruning strategy should achieve local optimization by preserving the current layer’s output, while also considering the global impact of pruning on subsequent layers. Firstly, we formulate a global objective function. Suppose token pruning is performed at layers $l_{1}<l_{2}<l_{3}$. For each pruned layer $l\in\{l_{1},l_{2},l_{3}\}$, we aim to select a subset of tokens $\mathbb{P}_{l}$ such that the difference between the pruned outputs $X_{pruned}^{l+1}$ and original outputs $X_{origin}^{l+1}$ is minimized. To quantify hidden vectors’ difference, we use a unified distance function $D(\cdot,\cdot)$ to measure the discrepancy between the outputs before and after pruning. Then our objective is to minimize the total output discrepancy across all pruned layers:
@@
Building upon the proposed local-global objective, we introduce our method. As shown in Figure 3, our approach divides token pruning into multiple stages denoted as $\mathcal{S}=\{s_{1},\dots,s_{n}\}$. Under a predefined pruning ratio $\alpha$, each stage retains a fixed fraction of image tokens from the previous stage. Since early pruning stages retain more tokens and influence the pruning decisions of later stages, their objectives need to emphasize token diversity. In contrast, deeper stages preserve fewer tokens and have less impact on subsequent stages. Therefore, we set the hyperparameter $\lambda_{i}$ to gradually increase across stages.
@@ 5 Experiment
All pruning experiments are conducted on 8 NVIDIA A800 GPUs using the HuggingFace Transformers library. To determine pruning stages, we randomly sample 64 instances from the LLaVA-655k liu2023improvedllava ; liu2024llavanext ; liu2023llava dataset and use the same set across all models and benchmarks, thus avoiding separate calibration for each benchmark. We gradually reduce the number of image tokens at each stage. In the early layers, we use a larger $\lambda$ value to focus more on global information, while in the deeper layers, we use a smaller lambda to emphasize local details. More implementation details for different models are provided in the see Appendix A.2. Similar to the implementation of PyramidDrop, we compute the required attention scores separately within the FlashAttn module at the specified pruning layers, achieving full compatibility with FlashAttn dao2023flashattention2 ; dao2022flashattention . It is worth noting that all our experiments are conducted with FlashAttention acceleration enabled.
@@ 5.1 Main results
As shown in Table 1, we conduct extensive experiments across different model families and parameter scales. Empirical results demonstrate that our approach consistently surpasses state-of-the-art methods on most benchmark tasks. Our method achieves 98% of the original average performance under a 22% compression rate across LLaVA models of different sizes. Moreover, our method consistently outperforms all models, achieving better results than both attention-based and diversity-based approaches. We also visualize the impact of different methods on layer outputs in Figure 5, our method preserves consistency with the original outputs at both local and global levels. The Appendix A.3 further provides visualizations of the spatial distribution of image tokens selected by various methods. Our method yields more effective token selection in deeper layers.
@@ 5.2 Efficiency analysis
FastV 860 0.37GB (66.6% $\downarrow$) 6.45 3.77s (11$\downarrow$) 92.6
DivPrune 633 0.28GB (74.7% $\downarrow$) 4.20 5.00s (17$\uparrow$) 99.1
BTP(ours) 633 0.28GB (74.7% $\downarrow$) 4.52 3.91s(7$\downarrow$) 98.9
@@ 5.3 Ablation study
Choice of balance factor value: We first analyze the effect of $\lambda$ in the local-global objective functions. This factor determines the trade-off at each layer between preserving local outputs and contributing to the global output. We gradually increase the value of $\lambda$. We define the ratio between the performance of the pruned model and that of the base model on the target task as the performance gain. The computation of performace performance gain is detailed in the Appendix A.4.
@@
As shown in Figure 6, we conduct experiments on both LLaVA-v1.5 and LLaVA-v1.6 to evaluate the effectiveness of our approach. As illustrated, model performance peaks when $\lambda$ balances the two objectives. In contrast, relying solely on either the local or global objective leads to a decline in performance. This highlights the importance of our method in effectively balancing the two objectives. Therefore, we use smaller $\lambda$ values (0.5–0.7) in the shallow layers and larger values (0.8–1.0) in the deeper layers.
Effectiveness of rebalanced attention and spatial diversity initialization: We then perform ablation studies on the attention rebalance module and the spatial initialization module. We experimented with various combinations of the two modules. The results are presented in Table 4. It can be observed that removing the attention rebalance module results in a significant degradation in model performance.
@@
0.131s 1478.1 57.3 84.4
This degradation arises from the inherent bias in attention mechanisms, where positional encodings tend to shift attention disproportionately toward later tokens, leading to suboptimal token selection. On the other hand, omitting the spatial initialization module causes a marked increase in inference latency, in some cases even surpassing that of the original unpruned model. This suggests that while pruning reduces token count, naive initialization can introduce computational overhead that negates the benefits of pruning, thereby limiting the method’s applicability in latency-sensitive real-world scenarios. This demonstrates the effectiveness of the proposed module in improving both model performance and inference speed.
@@
Effectiveness of calibration-based pruning stage selection: To evaluate the effectiveness of our proposed calibration-based pruning stage selection, we compare it with a baseline that uniformly divides the pruning stages according to the total number of decoder layers, under the same compression rate. Experimental results are shown in Table 5. We observe that our pruning layer selection method outperforms uniform selection. This is especially evident on Qwen2.5-VL, where uniform selection leads to a significant performance drop. We attribute this to differences in how Qwen2.5-VL processes image tokens as shown in Figure 4.

### v2

@@ Abstract
Large Vision-Language Models (LVLMs) have shown impressive performance across multi-modal tasks by encoding images into thousands of tokens. However, the large number of image tokens results in significant computational overhead, and the use of dynamic high-resolution inputs further increases this burden. Previous approaches have attempted to reduce the number of image tokens through token pruning, typically by selecting tokens based on attention scores or image token diversity. Through empirical studies, we observe that existing methods often overlook the joint impact of pruning on both the current layer’s output (local) and the outputs of subsequent layers (global), leading to suboptimal pruning decisions. To address this challenge, we propose Balanced Token Pruning (BTP), a plug-and-play method for pruning vision tokens. Specifically, our method utilizes a small calibration set to divide the pruning process into multiple stages. In the early stages, our method emphasizes the impact of pruning on subsequent layers, whereas in the deeper stages, the focus shifts toward preserving the consistency of local outputs. Extensive experiments across various LVLMs demonstrate the broad effectiveness of our approach on multiple benchmarks. Our method achieves a 78% compression rate while preserving 96.7% of the original models’ performance on average. Our code is available at https://github.com/EmbodiedCity/NeurIPS2025-Balanced-Token-Pruning.
@@ 1 Introduction
Recent advances in Large Vision-Language Models (LVLMs) chen2024internvl ; gao2024mini ; liu2023improvedllava ; liu2023llava ; wang2024qwen2 ; zha2025enablellm3dcapacity have substantially improved visual understanding. These models typically employ a visual encoder to convert images into discrete tokens, which are then processed jointly with textual tokens by a large language model backbone. The incorporation of visual information significantly increases the total number of input tokens bai2025qwen2 ; liu2024llavanext ; zhao2025embodiedr , a problem further amplified when handling high-resolution images. In edge applications such as emergency monitoring 10502270 ; wu2025monitorvlmavisionlanguageframework , logistics chen2024ddl ; zhang2025logisticsvlnvisionlanguagenavigationlowaltitude , and smart homes 10.1145/3706598.3713265 , models are typically deployed on devices like drones and unmanned vehicles jian2023pathgenerationwheeledrobots ; cui2024onboardvisionlanguagemodelspersonalized , which are constrained by limited memory and strict latency requirements. The excessive number of image tokens poses a major bottleneck for deployment, drawing increasing research interest in accelerating edge inference ruan2025edmambarethinkingefficientevent .
Prior studies alvar2025divprune have demonstrated that visual tokens often exhibit significant redundancy chen2024image ; lin2025boosting . Consequently, visual token pruning has been proposed as an effective strategy to reduce input redundancy and enhance computational efficiency yang2024visionzip ; shang2024llava ; huang2024dynamic ; zhang2025llava ; ye2024atp . Visual token pruning faces two fundamental challenges: identifying the most important visual tokens and determining the appropriate layers for pruning. Existing token pruning strategies can be broadly classified into two categories: attention-based methods that leverage text-image interactions chen2024image ; xing2024pyramiddrop ; meng2025plphpperlayerperheadvision , and diversity-based methods that exploit the heterogeneity of visual representations alvar2025divprune . However, the impact of their distinct optimization objectives on overall model performance remains underexplored, and a systematic comparison between them is largely absent. Moreover, when it comes to pruning layer selection, existing methods rely heavily on validation performance and manually defined settings, lacking principled guidance based on the model’s intrinsic properties.
@@
To address these problems, we first explore the nature of image token pruning from an intuitive perspective: its impact on the current layer’s (local) output and its influence on the outputs of subsequent pruning layers (global). We begin by visualizing the spatial distribution of image tokens that receive higher attention from text tokens across different layers. As shown in Figure 1, we observe that the image tokens attended by text tokens vary across different layers. This indicates that pruning solely based on the current layer tends to overlook its impact on subsequent layers. Then we further investigate the impact of different pruning methods on the model outputs. Specifically, we compare the hidden states of output tokens at different decoding positions under two pruning methods with those of the original model.
@@ 2.1 Large Vision-Language Models
Recent progress in large vision language models (LVLMs) has been substantially accelerated by the open-sourcing of foundation models like LLaMA touvron2023llama and Vicuna zheng2023judging . Representative models, including LLaVA liu2023improvedllava ; liu2024llavanext ; liu2023llava , Qwen-VL bai2025qwen2 ; wang2024qwen2 , and InternVL chen2024internvl ; gao2024mini leverage vision encoders radford2021learningtransferablevisualmodels ; li2022blipbootstrappinglanguageimagepretraining ; choraria2024semanticallygroundedqformerefficient to encode images into visual tokens, which are then integrated into the language model for unified multimodal representation and understanding gao2024embodiedcitybenchmarkplatformembodied . For example, LLaVA-1.5 encodes image into 576 visual tokens using a single-scale encoder. As these models increasingly support high-resolution visual inputs bai2025qwen2 ; liu2024llavanext , the number of visual tokens grows. Using a multi-resolution encoding strategy, LLaVA-NeXT can generate up to 2,880 tokens per image. Multimodal large models have been widely applied in various scenarios, including embodied agent 10966246 . The large number of image tokens limits its applicability in scenarios such as real-time applications ruan2025premamba4dstatespace .
@@ 2.2 Visual Token Pruning
Early efforts to reduce visual token redundancy primarily focus on attention-based pruning chen2024efficient ; huang2025dynamicllavaefficientmultimodallarge ; zhang2025llavaminiefficientimagevideo ; meng2025plphpperlayerperheadvision . For example, FastV chen2024image prunes visual tokens with low attention scores after the filtering layer, with subsequent layers processing only the remaining token. Another approach, VTW lin2025boosting , adopts a complete token elimination strategy, removing all visual tokens after a specified layer. PyramidDrop xing2024pyramiddrop introduces a more sophisticated approach, performing stage-wise pruning throughout the transformer, ranking visual tokens by their attention scores to the instruction token at each stage and progressively discarding the least informative ones. Compared to attention-based methods, diversity-based methods prioritize retaining a richer variety of semantic information. For instance, DivPrune alvar2025divprune formulate token pruning as a Max-Min Diversity Problem porumbel2011simple ; resende2010grasp . Additionally, some methods fuse remaining tokens into retained tokens through token fusion such as LLaVA-PruMerge shang2024llava and VisionZip yang2024visionzip . Different from prior methods, our method jointly considers the impact of pruning on both the current layer and subsequent layers.
@@ 3.1 Visual token processing
In the prefilling stage, images and texts are first encoded into embedding vectors (tokens), which are then processed by LVLM. We denote the input token sequence as $\mathbf{X}$ which consists of the system prompt $\mathbf{X}_{S}$, the image tokens $\mathbf{X}_{I}$ and text tokens $\mathbf{X}_{T}$, $\mathbf{X}$ = ($\mathbf{X}_{S}$$\mathbf{X}_{I}$$\mathbf{X}_{T}$) . $\mathbf{X}$ is then fed into the LLM backbone composed of N decoder layers. For the l-th decoder layer, we denote the input as $\mathbf{X}^{(l)}$ and the layer output $\mathbf{X}^{(l+1)}$ is:
@@ 4.1 Limitations of existing methods
We analyze the impact of pruning image tokens on the subsequent text and response tokens. From Equations 1 and 2, we can see that pruning image tokens at $l$-th layer mainly affects the layer output $\mathbf{X}^{(l+1)}$ by changing the attention output, which is a weighted sum of the value vectors $V_{l}$. If the norms of the $V_{l}$ are similar, selecting image tokens with high importance scores defined in 3 effectively reduces the difference between the layer output before and after pruning. We provide supporting evidence for this assumption in the Appendix 7.1. Formally, given original $l$-th layer output $\mathbf{X}_{origin}^{(l+1)}$ and pruned $l$-th layer output $\mathbf{X}^{(l+1)}_{pruned}$ , distance metric function $D(\cdot,\cdot)$, we can define the objective function $\mathcal{L}_{atten}$ of attention-based methods chen2024image ; xing2024pyramiddrop as:
@@ 4.2 Balanced token pruning with joint local and global objectives
Based on the above analysis, we argue that an effective token pruning strategy should achieve local optimization by preserving the current layer’s output, while also considering the global impact of pruning on subsequent layers. As shown in Equation 1, the model’s output depends on both the outputs of previous layers and the attention module of the current layer. Therefore, to ensure that the final output of the pruned model remains similar to that of the original model, we should maintain the similarity between the output of each pruned layer and its corresponding original output. Firstly, we formulate a global objective function. Suppose token pruning is performed at layers $l_{1}<l_{2}<l_{3}$. For each pruned layer $l\in\{l_{1},l_{2},l_{3}\}$, we aim to select a subset of tokens $\mathbb{P}_{l}$ such that the difference between the pruned outputs $X_{pruned}^{l+1}$ and original outputs $X_{origin}^{l+1}$ is minimized. To quantify hidden vectors’ difference, we use a unified distance function $D(\cdot,\cdot)$ to measure the discrepancy between the outputs before and after pruning. Then our objective is to minimize the total output discrepancy across all pruned layers:
@@
Building upon the proposed local-global objective, we introduce our method. Figure 3, our approach divides token pruning into multiple stages denoted as $\mathcal{S}=\{s_{1},\dots,s_{n}\}$. Under a predefined pruning ratio $\alpha$, each stage retains a fixed fraction of image tokens from the previous stage. As shown in Appendix 7.2, we can observe that retaining only a small number of image tokens is sufficient to optimize the attention objectives. Since early pruning stages retain more tokens and influence the pruning decisions of later stages, their objectives need to emphasize token diversity. In contrast, deeper stages preserve fewer tokens and have less impact on subsequent stages. Therefore, we set the hyperparameter $\lambda_{i}$ to gradually increase across stages.
@@ 5 Experiment
All pruning experiments are conducted on 8 NVIDIA A800 GPUs using the HuggingFace Transformers library. To determine pruning stages, we randomly sample 64 instances from the LLaVA-655k liu2023improvedllava ; liu2024llavanext ; liu2023llava dataset and use the same set across all models and benchmarks, thus avoiding separate calibration for each benchmark. We gradually reduce the number of image tokens at each stage. In the early layers, we use a larger $\lambda$ value to focus more on global information, while in the deeper layers, we use a smaller lambda to emphasize local details. More implementation details for different models are provided in the see Appendix 7.3. Similar to the implementation of PyramidDrop, we compute the required attention scores separately within the FlashAttn module at the specified pruning layers, achieving full compatibility with FlashAttn dao2023flashattention2 ; dao2022flashattention . It is worth noting that all our experiments are conducted with FlashAttention acceleration enabled.
@@ 5.1 Main results
As shown in Table 4.3, we conduct extensive experiments across different model families and parameter scales. Empirical results demonstrate that our approach consistently surpasses state-of-the-art methods on most benchmark tasks. Our method achieves 98% of the original average performance under a 22% compression rate across LLaVA models of different sizes. Moreover, our method consistently outperforms all models, achieving better results than both attention-based and diversity-based approaches. We also visualize the impact of different methods on layer outputs in Figure 5, our method preserves consistency with the original outputs at both local and global levels. The Appendix 7.5 further provides visualizations of the spatial distribution of image tokens selected by various methods. Our method yields more effective token selection in deeper layers.
@@ 5.2 Efficiency analysis
[…]
@@ 5.3 Ablation study
Choice of balance factor value: We first analyze the effect of $\lambda$ in the local-global objective functions. This factor determines the trade-off at each layer between preserving local outputs and contributing to the global output. To thoroughly analyze the contribution of each pruning layer, we perform comprehensive ablation experiments on the LLaVA model. Our method includes three pruning layers, and we evaluate three configurations by fixing the $\lambda$ parameters of two layers while varying the remaining one: (1) tuning the shallow layer while fixing the middle and deep layers, (2) tuning the middle layer while fixing the shallow and deep layers, and (3) tuning the deep layer while fixing the shallow and middle layers. We define the ratio between the performance of the pruned model and that of the base model on the target task as the performance gain. The computation of performace performance gain is detailed in the Appendix 7.4.
@@
As shown in Figure 6, we can observe an early preference for the diversity objective in the shallow layers results in performance degradation. The middle layers should still retain a moderate degree of diversity, whereas the deeper layers, due to the limited number of remaining tokens, should prioritize the attention objective. This highlights the importance of our method in effectively balancing the two objectives.
Effectiveness of rebalanced attention and spatial diversity initialization: We then perform ablation studies on the attention rebalance module and the spatial initialization module.
@@
0.231s 1478.1 57.3 84.4
We experimented with various combinations of the two modules. The results are presented in Table 4. It can be observed that removing the attention rebalance module results in a significant degradation in model performance. This degradation arises from the inherent bias in attention mechanisms, where positional encodings tend to shift attention disproportionately toward later tokens, leading to suboptimal token selection. On the other hand, omitting the spatial initialization module causes a marked increase in inference latency, in some cases even surpassing that of the original unpruned model. This suggests that while pruning reduces token count, naive initialization can introduce computational overhead that negates the benefits of pruning, thereby limiting the method’s applicability in latency-sensitive real-world scenarios zhang2025citynavagentaerialvisionandlanguagenavigation . This demonstrates the effectiveness of the proposed module in improving both model performance and inference speed. We also conducted an ablation study on the distance definitions used in the spatial diversity initialization module. As shown in the Appendix 7.7, we found that the Euclidean distance models the diversity of image tokens more effectively than the Manhattan distance.
Effectiveness of calibration-based pruning stage selection: To evaluate the effectiveness of our proposed calibration-based pruning stage selection, we compare it with a baseline that uniformly divides the pruning stages according to the total number of decoder layers, under the same compression rate. Experimental results are shown in Table 5.
@@
We observe that our pruning layer selection method outperforms uniform selection. This is especially evident on Qwen2.5-VL, where uniform selection leads to a significant performance drop. We attribute this to differences in how Qwen2.5-VL processes image tokens as shown in Figure 4. We also conduct an ablation study on the size and composition of the calibration set. Specifically, we expanded the calibration set by incorporating images from multiple datasets, including GQA, $V^{*}$Bench wu2023vguidedvisualsearch , and SQA and UrbanVideo-Bench zhao2025urbanvideobenchbenchmarkingvisionlanguagemodels . We then repeated the experiment shown in Figure 4 using calibration set sizes of 64, 128, and 256. The results are presented in Appendix 7.6, we can see that the variation patterns of image tokens remain consistent across different calibration set sizes and content, demonstrating the robustness of our pruning layer selection method.
Ablation on the Computation of Attention-Based Importance Scores: In our method, the importance score of each image token is obtained by using the attention assigned to it by the last text token in the prefilling stage. To verify the robustness of this design, we conduct an ablation study comparing different ways of computing the importance score: 1. Averaging attention weights from all text tokens to each image token. 2. Following the approach in zhang2025sparsevlmvisualtokensparsification , where image–text similarity is first computed and the most similar text tokens are then selected to calculate the importance score.
Table 6: Ablation study on importance score calculation method.
Method MME MMB POPE GQA SQA
last-token(ours) 1497 63.4 85.6 59.1 69.1
averaged-tokens 1490 62.8 84.7 57.3 69.4
similarity-based 1485 63.1 84.7 57.9 69.7
The results are shown in Table 6. We can see that last token efficiently modeling the importance score. We believe that the last token in the input prompt is a suitable choice for computing the importance score because it is typically decoded as the first output token during the decoding stage. This allows it to effectively capture the model’s focus.